OpenStudy (anonymous):

Differentiate y=e^x/x^2

OpenStudy (anonymous):

y'= ((xe^x) - (2e^x)) / x^3

OpenStudy (anonymous):

i got \[[(xe ^{x})(x+2)]/x ^{4}\]

OpenStudy (lgbasallote):

y' = [x^2(e^x) - e^x(2x)]/x^4 =[(e^x)(x^2-2x)]/x^4 =[x(e^x)(x-2)]/x^4 =[e^x(x-2)]/x^3

OpenStudy (anonymous):

oh yea...you can cross out one of the x's and yea you should get what lgsallote got